OpenStudy (anonymous):

\[\frac { \sqrt { 2 } -1 }{ \sqrt { 2 } +1 } =\quad a+b\sqrt { 2 } \quad then\quad find\quad the\quad value\quad of\quad a\quad and\quad b. \]

OpenStudy (anonymous):

Can ANyONe SOlVe It??

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

ok lets look at the left hand side first

OpenStudy (anonymous):

okay! carry on!!

OpenStudy (unklerhaukus):

\[LHS=\frac { \sqrt { 2 } -1 }{ \sqrt { 2 } +1 } \] \[=\frac { \sqrt { 2 } -1 }{ \sqrt { 2 } +1 } \times1\] \[=\frac { \sqrt { 2 } -1 }{ \sqrt { 2 } +1 } \times\frac{\sqrt 2-1}{\sqrt 2-1}\]

OpenStudy (unklerhaukus):

simplify

OpenStudy (anonymous):

hmm....

OpenStudy (anonymous):

Hey always if the denominator is a number, then it is for us to solve it. So first try to do that. If a square root exists in the denominator try to make it a number. Now think we've a sqrt her...so shd square it without changing the fraction......option is that use (a+b)(a-b)=aa-b2

OpenStudy (unklerhaukus):

\[=\frac {( \sqrt { 2 } -1)(\sqrt 2-1) }{ (\sqrt { 2 } +1 )(\sqrt 2-1)} \]

OpenStudy (anonymous):

say a+b is there in the denominator, multiply the fraction i.e both numerator and denom woth a-b. So in denom it'll be a2-b2......Solved

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hmmmmm...

OpenStudy (unklerhaukus):

what are you getting for the numerator and denominator ?

OpenStudy (anonymous):

m confused

OpenStudy (unklerhaukus):

\[( \sqrt { 2 } -1)(\sqrt 2-1) =\] \[{ (\sqrt { 2 } +1 )(\sqrt 2-1)}=\]

OpenStudy (anonymous):

A^2- b^2

OpenStudy (anonymous):

UNKLE GIVE the solution

OpenStudy (anonymous):

Thanks mathy

mathslover (mathslover):

got it ? ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

given is in the form (a-b)/(a+b)..... multiply with a-b to numerator & denom it becomes (a-b)2/(a2-b2)........read a2=Square of a substitute the values of a and b in this problem to get the answer

mathslover (mathslover):

welcome

OpenStudy (anonymous):

Indian gods

mathslover (mathslover):

Correct.... not Indian Gods but only Gods. These gods an worshipped all over world

OpenStudy (anonymous):

cool!!

OpenStudy (anonymous):

@mathslover how a= 3 &b = -2

mathslover (mathslover):

compare both eqn.

OpenStudy (anonymous):